The width of a rectangle is 12 less than its length. The area is 45 square inches. Find the dimensions of the rectangle.
Answer by TimothyLamb(4379)

w = L - 12
Lw = 45
---
L(L - 12) = 45
LL - 12L - 45 = 0
---
the above quadratic equation is in standard form, with a=1, b=-12, and c=-45
---
to solve the quadratic equation, by using the quadratic formula, copy and paste this:
1 -12 -45
into this solver: https://sooeet.com/math/quadratic-equation-solver.php
---
the two real roots (i.e. the two x-intercepts), of the quadratic are:
L = 15
L = -3
---
negative length doesn't make sense, so use the positive root
answer:
L = 15 in
w = 3 in
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php